OpenStudy (tester97):

Use the binomial theorem to expand the binomial. (s + 4v)^5 (1 point) here are my choices s5 – 20s4v + 160s3v2 – 640s2v3 + 640sv4 – 1024v5 s5 + 20s4v + 160s3 + 640s2+ 1280s + 1024 s5 + 20s4v + 160s3v2 + 640s2v3 + 1280sv4 + 1024v5 s5 + 20s4v + 80s3v2 + 640s2v3 + 640sv4 + 1024v5

OpenStudy (anonymous):

s^5 + 15s^4v + 90s^3v^2 + 270s^2v^3 + 405sv^4 + 243v^5

OpenStudy (tester97):

thats not a answer choice that i have

OpenStudy (anonymous):

http://www.purplemath.com/modules/binomial2.htm

OpenStudy (anonymous):

sorry i posted wrong 1. this link can help u

OpenStudy (anonymous):

I posted the link in chat.

OpenStudy (anonymous):

jenna its c

OpenStudy (wolfe8):

Tutorial: http://www.mathsisfun.com/algebra/binomial-theorem.html Expand the following: (s+4 v)^5 Expand (s+4 v)^5 using the binomial expansion theorem. (s+4 v)^5 = sum_(k=0)^5 binomial(5, k) (4 v)^(5-k) s^k = binomial(5, 0) (4 v)^5 s^0+binomial(5, 1) (4 v)^4 s^1+binomial(5, 2) (4 v)^3 s^2+binomial(5, 3) (4 v)^2 s^3+binomial(5, 4) (4 v)^1 s^4+binomial(5, 5) (4 v)^0 s^5: 1024 v^5 binomial(5, 0)+256 s v^4 binomial(5, 1)+64 s^2 v^3 binomial(5, 2)+16 s^3 v^2 binomial(5, 3)+4 s^4 v binomial(5, 4)+s^5 binomial(5, 5) Evaluate the binomial coefficients by looking at Pascal's triangle. binomial(5, 0) = 1, binomial(5, 1) = 5, binomial(5, 2) = 10, binomial(5, 3) = 10, binomial(5, 4) = 5 and binomial(5, 5) = 1: s^5+s^4×4 v×5+s^3 (4 v)^2 10+s^2 (4 v)^3 10+s (4 v)^4 5+(4 v)^5 Distribute exponents over products in (4 v)^2. Multiply each exponent in 4 v by 2: s^5+s^4×4 v×5+s^3×4^2 v^2 10+s^2 (4 v)^3 10+s (4 v)^4 5+(4 v)^5 Evaluate 4^2. 4^2 = 16: s^5+s^4×4 v×5+s^3×16 v^2 10+s^2 (4 v)^3 10+s (4 v)^4 5+(4 v)^5 Distribute exponents over products in (4 v)^3. Multiply each exponent in 4 v by 3: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×4^3 v^3 10+s (4 v)^4 5+(4 v)^5 In order to evaluate 4^3 express 4^3 as 4×4^2. 4^3 = 4×4^2: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×4×4^2 v^3 10+s (4 v)^4 5+(4 v)^5 Evaluate 4^2. 4^2 = 16: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×4×16 v^3 10+s (4 v)^4 5+(4 v)^5 Multiply 4 and 16 together. 4×16 = 64: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×64 v^3 10+s (4 v)^4 5+(4 v)^5 Distribute exponents over products in (4 v)^4. Multiply each exponent in 4 v by 4: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×64 v^3×10+s×4^4 v^4 5+(4 v)^5 Compute 4^4 by repeated squaring. 4^4 = (4^2)^2: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×64 v^3×10+s (4^2)^2 v^4 5+(4 v)^5 Evaluate 4^2. 4^2 = 16: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×64 v^3×10+s×16^2 v^4 5+(4 v)^5 Evaluate 16^2. | 1 | 6 × | 1 | 6 | 9 | 6 1 | 6 | 0 2 | 5 | 6: s^5+s^4×4 v×5+s^3×16 v^2×10+s^2×64 v^3×10+s×256 v^4 5+(4 v)^5 Multiply 4 and 5 together. 4×5 = 20: s^5+20 s^4 v+s^3×16 v^2×10+s^2×64 v^3×10+s×256 v^4×5+(4 v)^5 Multiply 16 and 10 together. 16×10 = 160: s^5+20 s^4 v+160 s^3 v^2+s^2×64 v^3×10+s×256 v^4×5+(4 v)^5 Multiply 64 and 10 together. 64×10 = 640: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+s×256 v^4×5+(4 v)^5 Multiply 256 and 5 together. 256×5 = 1280: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+(4 v)^5 Distribute exponents over products in (4 v)^5. Multiply each exponent in 4 v by 5: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+4^5 v^5 Compute 4^5 by repeated squaring. For example a^7 = a a^6 = a (a^3)^2 = a (a a^2)^2. 4^5 = 4×4^4 = 4 (4^2)^2: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+4 (4^2)^2 v^5 Evaluate 4^2. 4^2 = 16: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+4×16^2 v^5 Evaluate 16^2. | 1 | 6 × | 1 | 6 | 9 | 6 1 | 6 | 0 2 | 5 | 6: s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+4×256 v^5 Multiply 4 and 256 together. 4×256 = 1024: Answer: | | s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+1024 v^5

OpenStudy (anonymous):

The answer is s^5+20 s^4 v+160 s^3 v^2+640 s^2 v^3+1280 s v^4+1024 v^5

OpenStudy (tester97):

FINALLY!!!!!! i have the answer and the way to work it out thank you wolfe :)

OpenStudy (wolfe8):

xD You're welcome

OpenStudy (anonymous):

told ya it was c

OpenStudy (tester97):

it might take me a while to understand it all i will still be able to do so,

OpenStudy (anonymous):

lol tester. *gives tester medal*

OpenStudy (anonymous):

I called wolfe from chat ;)

OpenStudy (anonymous):

@wolfe8 please help the young lady

OpenStudy (wolfe8):

If I can pull off a stunt like that again, sure.

OpenStudy (anonymous):

lol. Tester wolfe i ready

OpenStudy (tester97):

ok here is the question give me a minute with the table Which of the following equations bestrepresents the regression line for the data given in the table above? y = x + 2 y = 2x – 2 y = –x – 2 y = x – 2

OpenStudy (wolfe8):

I want a new post and new medals. Lolz

OpenStudy (tester97):

 this isnt the best looking table but here ya go